OpenStudy (coconutjj):

Using Implicit Differentiation, take the derivative of d(x^2+yx^2+2(x^2)y+x^3)/dx

OpenStudy (coconutjj):

\[\frac{ d(x^2+yx^2+2x^2y+x^3) }{ dx }\]

OpenStudy (jhannybean):

\[d(x^2 +yx^2 +2(x^2)y+x^3)\]

OpenStudy (jhannybean):

Just take the derivative like you normally do, but wherever you see a y, take the derivative of the y function and tack on a y'

OpenStudy (coconutjj):

okay...

OpenStudy (jhannybean):

So what is the derivative of \(x^2\)?

OpenStudy (coconutjj):

2x

OpenStudy (jhannybean):

\(yx^2\)? Use product rule.

OpenStudy (coconutjj):

d(x^2)/dx y + dy/dx(x^2)

OpenStudy (coconutjj):

2xy + dy/dx(x^2)

OpenStudy (jhannybean):

Why do you write it like that, haha. I would simply write it as\[d(yx^2) = x^2y' +2xy\]

OpenStudy (jhannybean):

Yeah, same thing. Alright, good job \(\checkmark\)

OpenStudy (coconutjj):

mkay

OpenStudy (jhannybean):

Same thing with \(2x^2y\)

OpenStudy (coconutjj):

4xy + (2x^2)y(prime)

OpenStudy (jhannybean):

Oh and one more thing, is your function \(f(x) = x^2+yx^2+2(x^2)y+x^3\)?

OpenStudy (coconutjj):

yes

OpenStudy (jhannybean):

Okay, so far we have \[y' = 2x + x^2y'+2xy + 4xy+2x^2y'\]

OpenStudy (jhannybean):

Correct?

OpenStudy (coconutjj):

i believe so

OpenStudy (jhannybean):

Alright, now the derivative of \(x^3\)?

OpenStudy (coconutjj):

3x^2

OpenStudy (jhannybean):

And that finishes the derivative part.

OpenStudy (jhannybean):

SO we have \[y'= 2x + x^2y'+2xy + 4xy+2x^2y'+3x^2\]

OpenStudy (coconutjj):

yes

OpenStudy (jhannybean):

Now take everything that does not have a \(y'\) and subtract it so it's on the other side.

OpenStudy (jhannybean):

Take your \(y'\) on the left side, and subtract it so its on the right with all the rest of its friends.

OpenStudy (coconutjj):

Wait I think I get it.. you just factor out the d/dx right ?

OpenStudy (jhannybean):

yep.

OpenStudy (coconutjj):

alright thank-you... ill take from here

OpenStudy (jhannybean):

that's why I like writing it out as \(y'\) instead of \(\dfrac{dy}{dx}\)

OpenStudy (jhannybean):

Makes it look much prettier :)

OpenStudy (coconutjj):

THANK YOU~

OpenStudy (jhannybean):

No problem. Gl. happy New Year :O)